A gray kangaroo can bound across a flat stretch of
Answer #1

Range of gray kangaroo = 8.0 m

R=\frac{V^2\sin2\theta}{g}=8.0\,m

a) If angle of take off ,\theta = 180

Take off speed is V=\sqrt{8.0\frac{g}{\sin2\theta}}=11.6 \,m/s

b) Horizontal speed = v \cos\theta = 11 m/s